Miller Corporation acquired 30% of the outstanding common stock of Crowell Corporation for $160,000 on January 1, 2018, and obtained significant influence. The purchase price of the shares was equal to their book value. During 2018, the following information is available for Crowell:

Mar. 31 Declared and paid a cash dividend of $50,000.

June 30 Reported semiannual earnings of $120,000 for the first half of 2018.

Sept. 30 Declared and paid a cash dividend of $50,000.

Dec. 31 Reported semiannual earnings of $140,000 for the second half of 2018.

Required: 1. What is the balance in Miller’s investment account on December 31, 2018?

(It isn't $199,000 or $160,000)
Answer #1

Total net income of Crowell Corporation = 120,000 + 140,000

= $260,000

Total cash dividend paid by Crowell Corporation = 50,000 + 50,000

= $100,000

Cost of 30% common stock acquired of Crowell Corporation = $160,000

Revenue from investment in Crowell Corporation = Net income of Crowell Corporation x Percentage of common stock acquired

= 260,000 x 30%

= $78,000

Share of dividend from Crowell Corporation = Dividend paid by Crowell Corporation x Percentage of common stock acquired

= 100,000 x 30%

= $30,000

Balance in investment at December 31 = Cost of 25% common stock acquired of Crowell Corporation + Revenue from investment in Crowell Corporation - Share of dividend from Crowell Corporation

= 160,000 + 78,000 - 30,000

= $208,000

Balance in Miller’s investment account on December 31, 2018 = $208,000

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.